# reference.bib

### v1

@@ Abstract
Capturing high-quality visual content under low-light conditions remains a challenging problem due to severe noise, motion blur, and underexposure, which degrade the performance of downstream applications. Traditional frame-based low-light enhancement methods often amplify noise or fail to preserve structural details, especially in real-world scenarios. Event cameras, offering high dynamic range and microsecond temporal resolution by asynchronously capturing brightness changes, emerge as promising alternatives for low-light imaging. However, existing event-image fusion methods suffer from simplistic fusion strategies and inadequate handling of spatial-temporal misalignment and noise. To address these challenges, we propose EvRWKV, a novel framework that enables continuous cross-modal interaction through dual-domain processing. Our approach incorporates a Cross-RWKV module, leveraging the Receptance Weighted Key Value (RWKV) architecture for fine-grained temporal and cross-modal fusion, and an Event Image Spectral Fusion Enhancer (EISFE) module, which jointly performs adaptive frequency-domain noise suppression and spatial-domain deformable convolution alignment. Extensive qualitative and quantitative evaluations on real-world low-light datasets (SDE, SDSD, RELED) demonstrate that EvRWKV achieves state-of-the-art performance, effectively enhancing image quality by suppressing noise, restoring structural details, and improving visual clarity in challenging low-light conditions.
@@
Low-light image enhancement, event cameras, cross-modal fusion, RWKV, frequency-domain filtering, noise suppression, spatiotemporal alignment.
@@
I Introduction
Capturing high-quality visual content under low-light conditions remains a critical yet challenging task in computer vision [li2021low]. Images and videos acquired in such environments often suffer from severe noise, motion blur, and underexposure [li2021low], which degrade downstream applications such as autonomous driving and surveillance. Traditional frame-based low-light image enhancement methods [lamba2021restoring, wang2021seeing, fu2016weighted, wang2023low] have made progress, but often struggle to effectively address the inherent complexities of real-world low-light scenarios. Many methods [xu2022snr, cai2023retinexformer, wang2023ultra, 10810469, li2021low]can amplify existing noise while attempting to brighten the image, leading to visually unappealing results. Furthermore, approaches relying on synthetic data for training often fail to generalize well to the diverse and unpredictable nature of real-world low-light scenes.
Event cameras, using bio-inspired sensors [lichtsteiner2008128], are widely used in the field of computer vision due to their high dynamic range and microsecond level time resolution [scheerlinck2019ced, zheng2023deep], such as image restoration [jiang2023event], video super-resolution [lu2023learning], optical flow [luo2023learning], deblurring [jiang2020learning, kim2022event], etc., and have also become a promising alternative to low-light imaging scenes. By asynchronously capturing per-pixel brightness changes, events excel at preserving structural details even in extreme darkness and are less susceptible to motion blur. This makes them particularly attractive for low-light enhancement, where traditional cameras struggle due to signal-to-noise ratio limitations and motion artifacts. However, exploiting event cameras for low-light enhancement is not straightforward. Existing methods [stoffregen2020reducing, rebecq2019high] that rely solely on event cameras often face limitations due to the sparse nature of the event data, which can lead to poor spatial resolution and difficulty in handling noise under extreme low light conditions. Furthermore, fusion techniques combining events and images often use simplistic approaches [jiang2023event, wang2020event, liu2023low, liang2024towards], which fail to effectively capture the complex spatial-temporal interactions between them. As a result, these methods do not fully leverage the complementary strengths of event and image data, limiting their performance in low-light enhancement tasks.
A major limitation of current event and image fusion techniques lies in their simplistic handling of these distinct data streams. Many existing methods [jiang2023event, wang2020event, liu2023low] process events and images through separate pathways and then perform rudimentary fusion, typically by concatenating or adding features from both modalities. This decoupled approach fails to fully exploit the rich, cross-modal relationships inherent in simultaneously captured event and image data. In particular, fine-grained spatial-temporal interactions between the high-frequency dynamics of events and the textural information in images are often overlooked, leading to suboptimal fusion. Furthermore, both modalities are vulnerable to noise, including event noise in extreme darkness [gallego2020event] and image sensor noise [li2021low], and these basic fusion strategies may amplify such noise interference, resulting in further deterioration of enhancement quality. A fundamental challenge arises from the structural mismatch between the event data, which is temporally continuous but spatially sparse [lichtsteiner2008128], and the images, which offer dense spatial sampling but with limited temporal resolution [li2021low]. This discrepancy hinders effective integration of the motion cues from events with the structural details in images, preventing the achievement of optimal low-light image enhancement, as shown in Fig. 1 (c).
@@ I Introduction
To address these limitations, we introduce EvRWKV, a novel framework designed for continuous cross-modal interaction through dual-domain processing. Unlike earlier fusion strategies [jiang2023event, wang2020event, liu2023low], our EvRWKV achieves noise-robust spatiotemporal alignment through recursive dual-domain interaction and dynamic cross-modal gating, adaptively balancing event motion cues and image textures. It consists of two core modules: the Cross-RWKV module and the Event Image Spectral Fusion Enhancer (EISFE) module. The Cross-RWKV module is at the heart of our approach and pioneers the use of the Receptance Weighted Key Value (RWKV) architecture [peng2023rwkv] for cross-modal fusion in low-light enhancement. This module is designed to effectively capture temporal dependencies and enable fine-grained interaction between event and image features. Complementing this, EISFE addresses noise and structural degradation through coordinated frequency-spatial domain processing. In the frequency domain, EISFE employs an Adaptive Gaussian Filter to dynamically suppress noise while preserving structural information. Simultaneously, in the spatial domain, deformable convolutions with learnable offset fields actively align and fuse high-frequency details from both modalities, resolving spatial misalignment. By combining these two core modules, EvRWKV effectively enhances low-light images while addressing noise, structural degradation.
@@
In summary, EvRWKV represents a significant advancement in low-light image enhancement by effectively harnessing the synergistic potential of event and image cameras through a unified, cross-modal, and cross-domain architecture. The contributions of this work are as follows:
@@
  • We propose EvRWKV, a novel framework that establishes continuous cross-modal interaction through dual-domain processing for low-light enhancement, establishing intrinsic information exchange between event and image modalities.
@@
  • We demonstrate that EvRWKV achieves state-of-the-art (SOTA) performance on low-light enhancement across real-world datasets (SDE, SDSD, RELED), effectively suppressing noise and improving visual quality.
The rest of this paper is organized as follows: Section II details the related work in low image enhancement. Section III describes the workflow for each step of our EvRWKV method. Section IV gives a detailed overview of the experimental results and analysis. In the last section V, we summarize and look forward to our work.
@@ II Related Work
II-A Frame-based Low-light Enhancement
@@
Low-light enhancement(LLE) aims to improve the visibility and quality of images captured in dimly lit environments. Traditional frame-based low-light enhancement methods primarily operate on standard image frames. These methods can be broadly categorized into histogram equalization [arici2009histogram, abdullah2007dynamic, celik2011contextual], Gama Correction [huang2012efficient, rahman2016adaptive] and Retinex-based algorithms [guo2016lime, xu2020star, fu2016weighted, zhao2021retinexdip, liu2024efficient]. The first two methods directly enhance the intensity and contrast of low light image, while the Retinex-based algorithm models the image as a combination of illumination and reflectivity. Although effective in improving visibility, these approaches often encounter limitations in handling complex lighting conditions, especially in scenes with varying illumination or dark regions. Moreover, these methods may amplify noise, particularly in low-light regions, leading to undesirable artifacts.
@@
Recent advancements have been propelled by deep learning, particularly with convolutional neural networks (CNNs) [fu2016weighted, li2018structure, chen2018learning] and Transformer architectures [xu2022snr, cai2023retinexformer], which have led to the development of more sophisticated frame-based low-light enhancement methods. Chen et al. [chen2018learning] established a low light image dataset and proposed a full convolution network for enhancement. Wang et al. [wang2019underexposed] proposed to enhance underexposed photos by learning the illumination map. Wang et al. [wang2021seeing] collect SDSD dataset by using mechatronic system and proposed a framework integrating progressive alignment and Retinex-based illumination prediction. Cai et al. [cai2023retinexformer] propose a one-stage Transformer framework Retinexformer for low-light image enhancement, leveraging Retinex theory and illumination-guided attention to suppress noise. However, these frame-only methods fundamentally suffer from information loss in dark regions due to sensor noise floor limitations, leading to irreversible texture degradation and motion blur amplification.
@@
II-B Event-guided Low-light Enhancement
The emergence of event cameras has significantly revolutionized low-light perception by offering a new paradigm in image sensing. Unlike traditional frame-based cameras that capture images at fixed intervals, event cameras detect changes in the scene asynchronously, producing a continuous stream of events that encode pixel-level intensity changes with high temporal resolution. Rebecq et al. [rebecq2019high] propose a recurrent network to reconstruct high speed and high dynamic range videos from event camera data. Event-based approaches for LLE like NER-Net [liu2024seeing] leveraged event streams’ temporal continuity to address non-uniform illumination via learnable timestamp calibration. Subsequent works explored hybrid event-frame fusion strategies. Jiang et al. [jiang2023event] established joint feature learning through attention, while [liang2024towards] introduced the SDE dataset with precise spatial-temporal alignment and proposed SNR-guided feature selection in EvLight framework. Kim et al. [kim2024towards] propose an event-guided end-to-end framework for joint low-light video enhancement and deblurring on the real-world RELED dataset, utilizing temporal alignment and cross-modal spectral filtering. However, current fusion architectures are still limited by methods that rely on basic fusion strategies, which fail to effectively separate modality-specific noise patterns. Additionally, many of these approaches suffer from inflexible fusion operators that lack adaptive cross-modal interaction mechanisms. Furthermore, they often underutilize the event streams’ intrinsic motion information, hindering comprehensive dynamic scene modeling.
II-C Receptance Weighted Key Value
The Receptance Weighted Key Value (RWKV) architecture [peng2023rwkv] emerges as a transformative sequential modeling paradigm, synergizing the parallel computation advantages of transformers with the memory efficiency of recurrent networks. Departing from conventional attention mechanisms constrained by quadratic complexity, RWKV introduces time-shifted receptance gates and exponentially decaying key projections to maintain hidden states [peng2024eagle], enabling linear computational scaling with sequence length. In computer vision, RWKV has demonstrated remarkable versatility across diverse tasks [li2024survey]. For instance, Restore-RWKV [yang2024restore] employs an omnidirectional token translation layer and recurrent WKV attention to restore medical images [yang2025explicit] by capturing spatial dependencies, while PointRWKV [he2025pointrwkv] leverages linear complexity and multi-head matrix states to enhance geometric feature extraction in 3D point cloud processing. Vision-RWKV [duan2024vision] further extends its capabilities to high-resolution image analysis by reducing spatial aggregation complexity, outperforming traditional Vision Transformers (ViTs). In real-time video analytics, the TLS-RWKV framework [zhu2024tls] achieves efficient temporal pattern recognition for action detection. StyleRWKV [dai2024stylerwkv] utilizes RWKV models for efficient style transfer, enhancing global and local context while maintaining low memory usage and linear time complexity.
Despite these advances, RWKV’s application to cross-modal fusion scenarios—particularly for processing asynchronous, heterogeneous data streams like images and events—remains underdeveloped. Existing RWKV variants struggle with asynchronous temporal alignment and spatial-temporal sparsity gaps. While recent works like RWKV-CLIP [gu2024rwkv] explore vision-language alignment, they primarily focus on static multimodal interactions. Our work proposes a novel perspective: establishing continuous event-image synergy, where events resolve temporal ambiguities in blurry regions through high-frequency temporal updates, while images anchor sparse event data to structural priors via spatial-semantic grounding. This approach aligns with RWKV’s inherent strength in modeling long-range dependencies while mitigating its limitations in handling asynchronous modalities, opening new pathways for applications in motion deblurring, low-light enhancement, and neuromorphic vision systems.
@@ III Methodology
We propose EvRWKV, a novel framework designed for continuous cross-modal interaction through dual-domain processin that leverages the complementary advantages of low-light images and event streams to achieve robust illumination enhancement and noise suppression. As illustrated in Fig. 2, our framework comprises three key parts: 1) Feature Initialization, 2) Cross-modal Feature Restoration, 3) Frequency-aware Feature Fusion. The framework takes a low-light image $I$ and paired event stream $\{e_{k}\}_{k=1}^{N}$ as inputs, producing an enhanced image $I_{en}$. At first, a series of convolution layers extract features from low light images and event data, and then process them through the 4-level U-shaped encoder-decoder architecture composed of cross-RWKV composed of spatial and channel hybrid components. The output of the architecture passes the EISFE module is enhanced in frequency domain and spatial domain fusion processing to improve image quality.
@@ III-A Feature Initialization
Recent low-light enhancement methods demonstrate that preliminary enhancement facilitates subsequent restoration. Following Retinex theory [land1971lightness], an image $I\in\mathbb{R}^{H\times W\times 3}$ can be decomposed into reflectance $R\in\mathbb{R}^{H\times W\times 3}$ and illumination $L\in\mathbb{R}^{H\times W}$ as $I=R\odot L$, where $\odot$ denotes element-wise multiplication. Following [liang2024towards], we enhance the low-light image $I$ through:
@@
where $\hat{L}$ is the estimated illumination map, implementing rough enhancement while preserving original details. The event stream $\{e_{k}\}$ is converted into a voxel grid $E\in\mathbb{R}^{H\times W\times B}$ ($B=32$) to preserve spatio-temporal information. The image $I_{lu}$ and event voxel $E$ are both downsampled to $\frac{H}{2}\times\frac{W}{2}$ via a two-layer 3$\times$3 conv with strides 1 & 2, extracting features $X^{\text{img}},X^{\text{ev}}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$

### v2

# EvRWKV: A Continuous Interactive RWKV Framework for Effective Event-Guided Low-Light Image Enhancement

@@ Abstract
Capturing high-quality visual content under low-light conditions remains a challenging problem due to severe noise and underexposure, which degrade the performance of downstream applications. Traditional frame-based low-light image enhancement methods often amplify noise or fail to preserve structural details. Event cameras, offering high dynamic range and microsecond temporal resolution by asynchronously capturing brightness changes, emerge as a promising complement for low-light imaging. However, existing fusion methods fail to fully exploit this synergy, either by forcing modalities into a shared representation too early or by losing vital low-level correlations through isolated processing. To address these challenges, we propose EvRWKV, a novel framework that enables continuous cross-modal interaction through dual-domain processing. Our approach incorporates a Cross-RWKV module, leveraging the Receptance Weighted Key Value (RWKV) architecture for fine-grained temporal and cross-modal fusion, and an Event Image Spectral Fusion Enhancer (EISFE) module, which jointly performs adaptive frequency-domain noise suppression and spatial-domain deformable convolution alignment. This continuous interaction maintains feature consistency from low-level textures to high-level semantics. Extensive qualitative and quantitative evaluations on real-world low-light datasets (SDE, SDSD, RELED) demonstrate that EvRWKV achieves state-of-the-art performance, effectively enhancing image quality by suppressing noise, restoring structural details, and improving visual clarity in challenging low-light conditions.
@@
Low-light image enhancement, event cameras, cross-modal fusion, RWKV, dual-domain processing.
Figure 1: Overall framework of event-guided LLIE approaches. (a) Early fusion: combining image and event data at the input. (b) Late fusion: processing image and event data separately and merging features. (c) Continuous interaction fusion: enabling ongoing interaction between image and event data.
I Introduction
Capturing high-quality visual content under low-light conditions remains a critical yet challenging task in computer vision [26]. Images and videos acquired in such environments often suffer from severe noise, motion blur, and underexposure [26], which degrade downstream applications such as autonomous driving and surveillance. Traditional frame-based low-light image enhancement (LLIE) methods [23, 47, 9, 50] have made progress, but often struggle to effectively address the inherent complexities of real-world low-light scenarios. Many methods [55, 3, 49, 33, 27]can amplify existing noise while attempting to brighten the image, leading to visually unappealing results. Furthermore, approaches relying on synthetic data for training often fail to generalize well to the diverse and unpredictable nature of real-world low-light scenes.
Event cameras, using bio-inspired sensors [32], are widely used in the field of computer vision due to their high dynamic range and microsecond level time resolution [43, 59], such as image restoration [17], video super-resolution [37], optical flow [38], deblurring [18, 19, 6], etc., and have also become a promising alternative to low-light imaging scenes. By asynchronously capturing per-pixel brightness changes, events excel at preserving structural details even in extreme darkness and are less susceptible to motion blur. This makes them particularly attractive for low-light image enhancement, where traditional cameras struggle due to signal-to-noise ratio limitations and motion artifacts. However, exploiting event cameras for LLIE is not straightforward. Existing methods [44, 42] that rely solely on event cameras often face limitations due to the sparse nature of the event data, which can lead to poor spatial resolution and difficulty in handling noise under extreme low light conditions. Furthermore, fusion techniques combining events and images often use simplistic approaches [17, 46, 35, 30], which fail to effectively capture the complex spatial-temporal interactions between them. As a result, these methods do not fully leverage the complementary strengths of event and image data, limiting their performance in LLIE tasks.
@@ I Introduction
We argue that a core challenge lies in designing a fusion architecture that respects the profound heterogeneity of event and image modalities while enabling meaningful synergy between them. As shown in Fig. 1 (b), prevailing methods often adopt a decoupled design [30, 20, 45, 36, 17], processing event and image streams through separate, parallel pathways. This architectural choice inherently inhibits continuous interaction, restricting feature exchange to a single terminal fusion step. Consequently, vital low-level complementary correlations are lost during independent hierarchical processing. These include transient event edges corresponding to faint image textures. By the time the features are combined, they are too abstract to benefit from this lost complementary detail, leading to underutilized complementarity, semantic misalignment, and suboptimal enhancement, as shown in Fig. 2 (c).
Conversely, as shown in Fig. 1 (a), strategies [31, 35] that merge modalities at the input level struggle with a fundamental mismatch in data structure. Events are temporally dense but spatially sparse [32], whereas images offer dense spatial information, but are temporally discrete [26]. Forcing these inherently complementary representations into a common format at the outset creates a representational bottleneck, obscuring their distinct characteristics and preventing networks from effectively learning complementary relationships. This approach may also amplify noise from both sources [10, 26], degrading the fused representation before feature extraction begins. The fundamental limitation across these paradigms is the absence of a continuous adaptive interaction between the modalities throughout the feature learning process.
@@
To address this challenge, we introduce EvRWKV, a novel framework engineered for continuous cross-modal interaction through dual-domain processing. Instead of treating fusion as a single-stage process, EvRWKV establishes persistent cross-modal interactions between the RGB data and event streams at every network layer. This ensures that features from each modality dynamically inform and refine one another throughout the entire representation learning hierarchy. At the core of our approach, the Cross-RWKV module leverages the Receptance Weighted Key Value (RWKV) architecture [40] to enable fine-grained interaction between event and image features, capturing temporal dependencies and ensuring consistent spatiotemporal alignment. Complementing this, the Event Image Spectral Fusion Enhancer (EISFE) module addresses noise and structural degradation by applying an Adaptive Gaussian Filter in the frequency domain to suppress noise while preserving key structural details, and using deformable convolutions with learnable offset fields in the spatial domain to align and fuse high-frequency details from both modalities. By ensuring dynamic interaction across all stages of feature extraction—from low-level edges and textures, which are calibrated in real time, to mid-level structures that compensate for each other’s deficiencies, and high-level abstract features that are optimized for consistency—EvRWKV achieves seamless cross-modal fusion. This end-to-end collaboration maintains continuous feature alignment, preventing early information loss or late-stage disconnection, thus delivering robust LLIE that effectively preserves fine details, ensures global consistency, and suppresses noise and structural degradation.
@@
In summary, EvRWKV represents a significant advancement in LLIE by effectively harnessing the synergistic potential of event and image cameras through a unified, cross-modal, and cross-domain architecture. The contributions of this work are as follows:
@@
  • We propose EvRWKV, a novel framework that establishes continuous cross-modal interaction through dual-domain processing for LLIE, establishing intrinsic information exchange between event and image modalities.
@@
  • We demonstrate that EvRWKV achieves state-of-the-art (SOTA) performance on LLIE across real-world datasets (SDE, SDSD, RELED), effectively suppressing noise and improving visual quality.
@@ II Related Work
II-A Frame-based LLIE
@@
LLIE aims to improve the visibility and quality of images captured in dimly lit environments. Traditional frame-based LLIE methods primarily operate on standard image frames. These methods can be broadly categorized into histogram equalization [2, 1, 4], Gama Correction [16, 41] and Retinex-based algorithms [12, 54, 9, 58, 33]. The first two methods directly enhance the intensity and contrast of low light image, while the Retinex-based algorithm models the image as a combination of illumination and reflectivity. Although effective in improving visibility, these approaches often encounter limitations in handling complex lighting conditions, especially in scenes with varying illumination or dark regions. Moreover, these methods may amplify noise, particularly in low-light regions, leading to undesirable artifacts.
@@
Recent advancements have been propelled by deep learning, particularly with convolutional neural networks (CNNs) [9, 28, 5] and Transformer architectures [55, 3, 60], which have led to the development of more sophisticated frame-based LLIE methods. Chen et al. [5] established a low light image dataset and proposed a full convolution network for enhancement. Wang et al. [48] proposed to enhance underexposed photos by learning the illumination map. Wang et al. [47] collect SDSD dataset by using mechatronic system and proposed a framework integrating progressive alignment and Retinex-based illumination prediction. Cai et al. [3] propose a one-stage Transformer framework Retinexformer for LLIE, leveraging Retinex theory and illumination-guided attention to suppress noise. However, these frame-only methods fundamentally suffer from information loss in dark regions due to sensor noise floor limitations, leading to irreversible texture degradation and motion blur amplification.
II-B Event-guided LLIE
The emergence of event cameras has significantly revolutionized low-light perception by offering a new paradigm in image sensing. Unlike traditional frame-based cameras that capture images at fixed intervals, event cameras detect changes in the scene asynchronously, producing a continuous stream of events that encode pixel-level intensity changes with high temporal resolution. Rebecq et al. [42] propose a recurrent network to reconstruct high speed and high dynamic range videos from event camera data. Event-based approaches for LLE like NER-Net [34] leveraged event streams’ temporal continuity to address non-uniform illumination via learnable timestamp calibration. Subsequent works explored hybrid event-frame fusion strategies. Jiang et al. [17] established joint feature learning through attention, while [30] introduced the SDE dataset with precise spatial-temporal alignment and proposed SNR-guided feature selection in EvLight framework. Kim et al. [20] propose an event-guided end-to-end framework for joint low-light video enhancement and deblurring on the real-world RELED dataset, utilizing temporal alignment and cross-modal spectral filtering. However, current fusion architectures are still limited by methods that rely on basic fusion strategies, which fail to effectively separate modality-specific noise patterns. Additionally, many of these approaches suffer from inflexible fusion operators that lack adaptive cross-modal interaction mechanisms. Furthermore, they often underutilize the event streams’ intrinsic motion information, hindering comprehensive dynamic scene modeling.
II-C Receptance Weighted Key Value
The Receptance Weighted Key Value (RWKV) architecture [40] emerges as a transformative sequential modeling paradigm, synergizing the parallel computation advantages of transformers with the memory efficiency of recurrent networks. Departing from conventional attention mechanisms constrained by quadratic complexity, RWKV introduces time-shifted receptance gates and exponentially decaying key projections to maintain hidden states [39], enabling linear computational scaling with sequence length. In computer vision, RWKV has demonstrated remarkable versatility across diverse tasks [29]. For instance, Restore-RWKV [57] employs an omnidirectional token translation layer and recurrent WKV attention to restore medical images [56] by capturing spatial dependencies, while PointRWKV [13] leverages linear complexity and multi-head matrix states to enhance geometric feature extraction in 3D point cloud processing. Vision-RWKV [8] further extends its capabilities to high-resolution image analysis by reducing spatial aggregation complexity, outperforming traditional Vision Transformers (ViTs). In real-time video analytics, the TLS-RWKV framework [61] achieves efficient temporal pattern recognition for action detection. StyleRWKV [7] utilizes RWKV models for efficient style transfer, enhancing global and local context while maintaining low memory usage and linear time complexity.
Despite these advances, RWKV’s application to cross-modal fusion scenarios—particularly for processing asynchronous, heterogeneous data streams like images and events—remains underdeveloped. Existing RWKV variants struggle with asynchronous temporal alignment and spatial-temporal sparsity gaps. While recent works like RWKV-CLIP [11] explore vision-language alignment, they primarily focus on static multimodal interactions. Our work proposes a novel perspective: establishing continuous event-image synergy, where events resolve temporal ambiguities in blurry regions through high-frequency temporal updates, while images anchor sparse event data to structural priors via spatial-semantic grounding. This approach aligns with RWKV’s inherent strength in modeling long-range dependencies while mitigating its limitations in handling asynchronous modalities, opening new pathways for applications in motion deblurring, LLIE, and neuromorphic vision systems.
@@ III Methodology
We propose EvRWKV, a novel framework designed for continuous cross-modal interaction through dual-domain processin that leverages the complementary advantages of RGB data and event streams to achieve robust illumination enhancement and noise suppression. As illustrated in Fig. 3, our framework comprises three key parts: 1) Feature Initialization, 2) Cross-modal Feature Restoration, 3) Frequency-aware Feature Fusion. The framework takes a low-light image $I$ and paired event stream $\{e_{k}\}_{k=1}^{N}$ as inputs, producing an enhanced image $I_{en}$. At first, a series of convolution layers extract features from low light images and event data, and then process them through the 4-level U-shaped encoder-decoder architecture composed of cross-RWKV composed of spatial and channel hybrid components. The output of the architecture passes the EISFE module is enhanced in frequency domain and spatial domain fusion processing to improve image quality.
@@ III-A Feature Initialization
Recent LLIE methods demonstrate that preliminary enhancement facilitates subsequent restoration. Following Retinex theory [24], an image $I\in\mathbb{R}^{H\times W\times 3}$ can be decomposed into reflectance $R\in\mathbb{R}^{H\times W\times 3}$ and illumination $L\in\mathbb{R}^{H\times W}$ as $I=R\odot L$, where $\odot$ denotes element-wise multiplication. Following EvLight [30], we enhance the low-light image $I$ through:
@@
where $\hat{L}$ is the estimated illumination map, implementing rough enhancement while preserving original details. The event stream $\{e_{k}\}$ is converted into a voxel grid $E\in\mathbb{R}^{H\times W\times B}$ ($B=32$) to preserve spatio-temporal information. The image $I_{lu}$ and event voxel $E$ are both downsampled to $\frac{H}{2}\times\frac{W}{2}$ via a two-layer $\times3$ conv with strides 1 & 2, extracting features $X^{\text{img}},X^{\text{ev}}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$ .